A chemistry teacher needs to strengthen a 12% solution of acid to strength s by mixing it with pure acid. How much pure acid must be added to 100ml of the 12% solution to make acid strength s?

Answer by jorel1380(3719)

100(.12)+x=s(100+x)
12+x=100s+sx
x-sx=100s-12
x(1-s)=100s-12
x=100s-12/1-s..